### v1

# HP-BERT: A framework for longitudinal study of Hinduphobia on social media via LLMs

Abstract
During the COVID-19 pandemic, community tensions intensified, fuelling Hinduphobic sentiments and discrimination against individuals of Hindu descent within India and worldwide. Large language models (LLMs) have become prominent in natural language processing (NLP) tasks and social media analysis, enabling longitudinal studies of platforms like X (formerly Twitter) for specific issues during COVID-19. We present an abuse detection and sentiment analysis framework that offers a longitudinal analysis of Hinduphobia on X (Twitter) during and after the COVID-19 pandemic. This framework assesses the prevalence and intensity of Hinduphobic discourse, capturing elements such as derogatory jokes and racist remarks through sentiment analysis and abuse detection from pre-trained and fine-tuned LLMs. Additionally, we curate and publish a ”Hinduphobic COVID-19 X (Twitter) Dataset” of 8,000 tweets annotated for Hinduphobic abuse detection, which is used to fine-tune a BERT model, resulting in the development of the Hinduphobic BERT (HP-BERT) model. We then further fine-tune HP-BERT using the SenWave dataset for multi-label sentiment analysis. Our study encompasses approximately 27.4 million tweets from six countries, including Australia, Brazil, India, Indonesia, Japan, and the United Kingdom. Our findings reveal a strong correlation between spikes in COVID-19 cases and surges in Hinduphobic rhetoric, highlighting how political narratives, misinformation, and targeted jokes contributed to communal polarisation. These insights provide valuable guidance for developing strategies to mitigate communal tensions in future crises, both locally and globally. We advocate implementing automated monitoring and removal of such content on social media to curb divisive discourse.
keywords:
Hinduphobia , COVID-19 , X (Twitter) , Large Language Models
journal: Nuclear
[…]
[inst1]organization=Department of Data Science and Artificial Intelligence,addressline=IIIT, Naya Raipur, postcode=Chhattisgarh, country=India
[…]
[inst2]organization=Transitional Artificial Intelligence Research Group, School of Mathematics and Statistics, addressline=UNSW Sydney, postcode=Sydney, country=Australia
[…]
[inst3]organization=Centre for Hindu Studies, addressline=Pingala Institute, postcode=Sydney, country=Australia
1 Introduction
A global pandemic emerged from the severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2) [1], which resulted in the coronavirus disease 2019 (COVID-19)[2] with a profound impact on social and economic activities worldwide [3, 4, 5]. One of the most significant aspects of this pandemic has been how social media has influenced public attitudes and behaviours related to the pandemic [6]. The COVID-19 pandemic sparked a global public health emergency and led to a rise in online discussions that exacerbated existing social divisions and sparked new conflicts within communities worldwide [7]. In India’s case, a country known for its diverse cultural landscape and inclusive principles, the pandemic exacerbated existing socio-political divisions, particularly concerning communal interactions [8]. Although efforts have focused on addressing anti-Islamic sentiments and behaviours [9, 10, 11, 12], there is another significant but less explored issue: the escalation of Hinduphobia [13, 14] during the COVID-19 crisis. Furthermore, the COVID-19 pandemic has also brought to light the issue of Sinophobia [15, 16, 17], but there has been little attention given by researchers to the rise of Hinduphobia.
Hinduism is the world’s oldest practising religion [18] with a population of more than 1.1 billion followers worldwide who mostly reside in India. Hinduism encompasses some of the oldest philosophical (Upanishads and Bhagavad Gita) [19, 20], epic and literary texts (Ramayana and Mahabharata) [21, 22]. Hinduism has contributed to modern science and technology with the invention of the binary and decimal number system [23] (also known as Hindu and Arabic numerals) [24] and systematic approach to scientific investigation using the philosophy of pramana (evidence) [25, 26]. Hinduism is based on non-violent principles that promote vegetarianism [27], with India having the largest vegetarian population in the world with the majority of Hindus being partially vegetarian (specific days of vegetarian diet) [27]. Hinduism urges taking a violent stand in self-defence [28] and the fight for dharma (ethics) as promoted in the Bhagavad Gita [29]. The term Hinduphobia refers to hate speech and racial and cultural slurs about Hindu religious and cultural practises [13]. Although limited, few studies have investigated the nature of Hinduphobia in television and popular culture [13], which became more prominent with the rise of the Internet and social media [14].
Hinduism in India withstood its grounds with a history of violence and discrimination given 800 years of foreign rule, with invasions leading to temple desecration [30]. Thousands of Hindu temples have been destroyed desecrated and converted into mosques and churches in history due to foreign rule with religious and political intervention [30, 31]. This has been continuing with Hinduphobia in social media and political discourses, especially in countries where Hindus have been a minority, such as Pakistan and Bangladesh [32, 33]. In a majority Hindu India, Hindupbois and anti-Hindu political rhetoric led to the genocide and exodus of Hindus [34] from the Kashmir valley region in the 1980s [35]. Therefore, Hinduphobia is not a recent phenomenon, it has more than a thousand years of history given the invasions and foreign rule of India [36]. Furthermore, Hindupbobic rhetoric has been one of the reasons for forces conversion of Hindu women in Pakistan [37] and gradual changes in ownership and conversion of Hindu temples [38]. This has been an issue not just in India but in countries with a Hindu history such as Indonesia [39].
The Kumbh Mela which is a major Hindu festival in India was associated with a high number of COVID-19 cases during the Delta variant outbreak in April 2022 [40, 41, 42]. This association, driven by a political media narrative, led to negative portrayals of the Hindu community on social media platforms including Twitter, Reddit, Instagram, and Facebook. These negative portrayals resulted in trends with hashtags such as #HinduVirus and #CoronaHindus, contributing to Hindubhobic statements and hate speech against the Hindu community. Previous studies have examined Hinduphobia in contexts of historical events, political dynamics, academic biases, interreligious relations, diaspora experiences, and cultural practices [13, 43]. Furthermore, there has been an upsurge in hate crimes against Hindus, as demonstrated by the cases of Khaniya Lal, Umesh Kohle, and Nupur Sharma [44]. These incidents highlight the escalating hostility towards Hindus during the pandemic and the need for religious tolerance in society. The COVID-19 pandemic has exacerbated communal tensions in India, which motivates our study to systematically study the rise of Hinduphobic sentiments connected to COVID-19 in India and the rest of the world.
In recent years, LLMs have emerged as powerful tools for analysing and understanding public sentiment on social media. LLMs, such as (GPT-3, PaLM, and GPT-4) [45, 46] can process vast amounts of textual data, revealing underlying patterns, sentiments, and discourse dynamics that shape public perception and societal narratives [47]. These models can be based on pre-training from a large data corpus and achieve strong performance accuracy in tasks such as decision-making, predictions, text generation and text summarizations [48]. LLMs can detect nuanced language patterns and emerging trends, enabling researchers and policymakers to respond more effectively to rising social and political tensions [49] taking into account security and privacy [48]. Addressing the root causes of discrimination and fostering social cohesion is crucial; however, applying LLMs in this context has a lot of challenges. Data privacy, ethical considerations, and the potential for algorithmic bias must be carefully managed to ensure the deployment of LLMs for societal well-being [50]. LLM must limit biases and maintain fairness, particularly when dealing with sensitive topics such as communal sentiments and hate speech [51].
Sentiment and semantic analysis have become crucial components of NLP applications, particularly for analysing vast amounts of unstructured textual data from social media platforms [52, 53]. These methods involve understanding the emotional tone (sentiment) and the meaning (semantics) of textual content to extract insights regarding public opinion, trends, and biases [54]. LLMs have significantly improved the accuracy and scalability of these tasks by leveraging massive datasets and deep model architectures. LLMs can capture complex linguistic patterns and contextual nuances previously challenging to NLP. One of the earliest LLMs, the Bidirectional Encoder Representations from Transformers (BERT) model [55], has been widely adopted for sentiment analysis tasks [56]. Several customised BERT models have been developed, including news recommender BERT [57], fake news BERT [58], legal-BERT [59, 60] and medical-BERT [61] for respective domains. The Multilingual-BERT (mBERT) [62] has been pre-trained on monolingual corpora in 104 languages, which has been designed to detect hate speech across multiple languages. DistilBERT [63] is a variant of BERT optimised for resource-limited applications. Moreover, advanced variants such as Ro-BERT-a [64] have proven effective for nuanced hate speech detection across diverse settings. In the case of detection of abuse and violence in NLP applications, BERT has been designed for various contexts and categories of hate speech. HateBERT [65] is trained on RAL-E, a large-scale dataset of Reddit comments from communities banned for being offensive, abusive, and hateful. These examples of customised BERT models and their variants demonstrate the adaptability and power of LLMs in addressing the complexities of hate speech detection in specific contexts, showcasing their impact in real-world NLP applications.
Sentiment and semantic analysis have been used to study various domains using LLMs, including social media and text analysis. Wang and Chandra [17] used BERT-based sentiment analysis and reported that surges in COVID-19 cases correlated with spikes in Sinophobic sentiments. Similarly, Chandra et al. [66] used BERT-based sentiment analysis to study trends in The Guardian newspaper coverage of COVID-19, revealing a predominance of negative sentiments. In another study, Chandra and Kulkarni [67] utilised semantic and sentiment analysis to evaluate the Bhagavad Gita Sanskrit-English translations, comparing verse-by-verse consistency across selected translators. In the field of machine translation, Wang et al. [68] evaluated Google Translate for its accuracy in translating Mandarin Chinese, applying sentiment and semantic analysis to assess translation precision. Similarly, Shukla et al. [69] analysed Sanskrit-to-English translations of the Bhagavad Gita, uncovering inconsistencies in translating philosophical and metaphorical terms of translation by Google Translate. Furthermore, Chandra and Saini [70] modelled the United States 2020 presidential elections using BERT-based sentiment analysis, providing insights into sentiment trends on social media before the elections. Chandra and Krishna [71] compared LSTM and BERT models to track public sentiment during the rise of COVID-19 cases in India. These studies collectively highlight the role of BERT-based models and LLMs in analysing sentiments across social and news media, illustrating their versatility and potential for understanding public opinion.
Furthermore, by leveraging the capabilities of LLMs, it is possible to conduct a more granular analysis of sentiment dynamics across different regions, demographics, and periods. Although we can more easily detect sentiments in expressions, the associated meaning is difficult to decipher for models that are not aware of the context and historical development, such as the case of Xenophobic [72, 73], Sinophobic [17] and Hinduphobic remarks. This challenge has been demonstrated for code-mixed language settings, such as the case of Hinglish using BERT [74] for hate-speech detection based on a database of Hinglish for misogyny and aggression [75]. Finally, in the case of Hinduphobic remarks, we do not have pre-existing labelled datasets to refine BERT-based models.
In this study, we present the ”Hinduphobic COVID-19 X (Twitter) Dataset, the first dataset of its kind, with over 8,000 tweets collected from November 2019 to December 2022 during the COVID-19 pandemic. It has 2000 manually labelled tweets and the rest employs GPT-3.5 turbo API. The dataset is labelled into two categories: Hinduphobic (negative) and positive/neutral, with all tweets related to Hinduism. This dataset is publicly available via Kaggle [^1] https://www.kaggle.com/datasets/ashutoshsingh22102/hinduphobic-covid-19-x-twitter-dataset-india. We use the Hinduphobia dataset to fine-tune a BERT (HP-BERT) model for classifying Hinduphobic tweets. Through a multi-stage fine-tuning strategy [76, 77, 78], we further train this model on the ”SenWave dataset” [79] to enable sentiment analysis. Subsequently, the model is applied to the ”Global COVID-19 Twitter dataset” [80], which includes tweets from six countries including Australia, Brazil, India, Indonesia, Japan, and the United Kingdom capturing a longitudinal perspective on Hinduphobic sentiment trends across these regions during and post-COVID-19 pandemic. We also analyze sentiment polarity scores to gain deeper insights into how Hinduphobic sentiments evolved and varied geographically. Additionally, we use HateBERT [65] to examine toxicity and abusive language in tweets in the ”Global COVID-19 X (Twitter) Dataset” [80], offering a further understanding of the propagation of Hinduphobia and the behavioural patterns of users spreading such content. We present a refined BERT (HP-BERT) model specifically designed for detecting Hinduphobia, with adaptations to handle Hinglish (Hindi and English) data from India [81], which includes content compiled for abuse and misogyny analysis [75]. This model effectively detects Hinduphobic content and performs sentiment analysis on identified tweets, providing nuanced insights into the emotional tone and context surrounding Hinduphobic discourse.
The rest of the paper is organised as follows: Section 2 reviews the literature relevant to this study. Section 3 describes the datasets used, including the data collection and preprocessing processes, along with details of the training and testing of our BERT (HP-BERT) model. Section 4 introduces the framework for our approach. Section 5 presents the study’s findings, discussing the performance of the HP-BERT model. Finally, Section 6 provides key insights, and Section 7 concludes the paper with future research directions.
2 Related Work
The study of online activity drew significant attention from the research community due to the unique set of challenges faced during the COVID-19 pandemic, with concerns such as anti-vaccination and misinformation [82, 83, 84]. The substantially increased social media activity provided researchers with a wealth of data to analyze public sentiment and discourse. This surge in online interactions also exacerbated stress in the society [85, 86], as individuals grappled with the rapid spread of misinformation, social isolation, and heightened fears about the pandemic. Chen et al. [87] curated a multilingual coronavirus Twitter dataset with over 123 million tweets for analysing the pandemic’s online discourse, reflecting major events and their impacts. Tahmasbi et al. [88] conducted one of the first studies on Sinophobic behaviour during the pandemic, analysing over 222 million tweets to reveal the rise in anti-Chinese sentiment.
Cho et al. [89] focused on privacy concerns related to COVID-19 contact tracing apps such as TraceTogether, exploring cryptographic protocols and privacy-preserving techniques. The study highlighted the balance between privacy protection and public health benefits, emphasising effective anonymisation and additional privacy measures. Ferrara et al. [90] investigated the influence of X (Twitter) bots during the early COVID-19 outbreak, particularly in spreading political conspiracies and misinformation. Their curated dataset revealed how bots amplified divisive content during global emergencies. Kouzy et al. [91] examined COVID-19 misinformation on Twitter, identifying its rapid spread and sources. They found misinformation often originated from informal groups rather than verified accounts, underscoring the need for robust monitoring and fact-checking mechanisms. Tahmasbi et al. [88] analysed Sinophobic content on social media platforms post-COVID-19, using word2vec models to trace the evolution of anti-Chinese sentiment. They highlighted the exacerbation of racial hostility during the pandemic across different online forums.
He et al. [92] developed a text classifier to detect and mitigate anti-Asian hate speech on Twitter during COVID-19, exploring counter-speech strategies to combat online racism effectively. Chandra et al. [93] employed LSTM models to forecast COVID-19 infections across Indian states, demonstrating their efficacy in predicting cases and providing a two-month-ahead prediction using a recursive prediction strategy. They also highlighted COVID-19’s broader impacts, addressing challenges in data reliability and population dynamics. Lande et al. [94] applied deep learning techniques for COVID-19 topic modelling for India on Twitter (X) using BERT-based approaches to identify recurring themes such as governance and vaccination. Their study correlated these topics with media coverage across different pandemic phases, providing insights into public discourse and information dissemination during crises.
Figure 1: Data collection and processing workflow from initial data extraction via Twitter API to manually labelling data points. This is followed by automatic labelling of the remaining tweets using GPT-3.5 Turbo using human-in-the-loop strategy [95].
Despite the extensive research on various aspects of the COVID-19 pandemic, there remains a gap in quantitative analysis regarding Hinduphobic sentiments during this period. This paper aims to fill this gap by examining the prevalence and impact of Hinduphobic discourse on social media platforms amidst the COVID-19 pandemic.
3 Methods
3.1 Datasets
In this study, we utilise three datasets that include two existing datasets and one curated by us:
  1. 1.
    We curate a new dataset by extracting and processing the Hinduphobic COVID-19 X (Twitter) dataset for the classification of Hinduphobic content, with fine-tuning to allow the HP-BERT model to learn the language and sentiment patterns specific to Hinduphobic tweets.
  2. 2.
    We use the SenWave dataset to fine-tune the HP-BERT model for Hinduphobic sentiment analysis. This multi-stage fine-tuning process helped the model adapt to sentiment classification tasks while retaining the domain-specific features learned from the first stage.
  3. 3.
    We employ the Global COVID-19 X (Twitter) dataset for sentiment analysis on Hinduphobic content, enabling the review of sentiment in tweets related to Hinduphobia globally during COVID-19.
3.1.1 Hinduphobic COVID-19 X (Twitter) Dataset
We extracted the Hinduphobic tweets and curated a dataset using the X (Twitter) application programmer interface (API) [^2] https://developer.x.com/en/docs/x-api. The dataset spans from April 2020 to January 2024, capturing the context of the COVID-19 pandemic with tweets originating from India, featuring both Hinglish (a mix of Hindi and English) and English. Each tweet includes terms such as COVID, COVID-19, or coronavirus, establishing a clear relation between the pandemic and the Hindu community. We employed targeted keywords and hashtags to identify negative (Hinduphobic) and positive/neutral sentiments associated with Hinduism. We present some of the terms identified and their relevance to Hinduphobic sentiments in (Table 1) along with positive/neutral terms. The negative terms such as ”Hindu superspreaders” (referring to the Kumbh Mela Hindu festival), ”Hindu rituals causing COVID” (related to temple gatherings), and derogatory references such as ”piss drinker” (associated with cow urine and dung) were used to malign and blame the Hindu community for the spread of COVID-19. Notable hashtags promoting Hinduphobic content included #CoronaJihad, #HindusSpreadCorona, #HinduCOVIDConspiracy, #COVID19HinduHate, and #COVIDHindutva. For the positive or neutral sentiment, we focused on terms associated with Hindu contributions to COVID-19 relief efforts, such as ”Namaste” or ”Namaskar” (a Hindu greeting that avoids physical contact) and references to organisations such as ISKCON (International Society of Krishna Consciousness), RSS (Rashtriya Swayamsevak Sangh), Swaminarayan Temple, and Mahavir Temple. These organisations played significant roles in COVID-19 relief by providing food, plasma donations, hospital support, and financial aid. Positive hashtags in this category included #NamasteForSafety, #HindusForHumanity, #ISKCONRelief, #RSSHelpingHands, #TempleAidCOVID, and #HinduCOVIDRelief. We derived the full list of keywords from scholarly sources [96], Hindu American Foundation [^3] https://www.hinduamerican.org/hindu-hate-glossary, news articles [^4] https://shorturl.at/3Dvrl, and Wikipedia [^5] https://en.wikipedia.org/wiki/Anti-Hindu_sentiment with explanations for each term’s association with negative or positive sentiments available in a document [^6] https://github.com/pinglainstitute/Hinduphobia-COVID-19-beyond.
Keyword Negative (Hinduphobic) Comments
Cow Urine, You’re right. Hindus nearly drink cow urine and bathe in cow dung. I saw videos during COVID suggesting bathing in cow dung as a way to prevent the virus. Disgusting folks.
Hindu Rituals People are performing corona puja in Bihar, West Bengal, treating coronavirus as a goddess and offering 9 laddus, 9 cloves, and 9 flowers so the goddess will leave and people won’t get sick. No masks, no social distancing. Will superstition ever die? Will scientific temper ever develop?
Temple Right after #AYODHYAVERDICT, the anti-Hindu crowd said, We need world-class hospitals, not temples, but Hindu temples have become resources in this crisis, building COVID hospitals. #HinduTemplesWithNation
Kumbh Why blame the Kumbh just because it’s a Hindu festival? Some new research in The Lancet may show that COVID doesn’t spread at farmer rallies or religious gatherings of other religions. The judiciary also seems to agree.
Festivals BBC News - India Covid: First UK aid arrives as coronavirus deaths mount. Modi and the BJP must take full blame for this tsunami of COVID-19 infections for his irresponsible opening of Hindu festivals that led to this disaster.
Keyword Positive/Neutral (Pro-Hindu) Comments
Namaste For the last few months, we all are fighting against COVID-19. Today, people all over the world are following Hindu practices such as doing Namaste instead of shaking hands, cremating bodies of COVID-19 casualties, etc., to stay safe.
Food Distribution It’s been 45 days since Hindu groups started distributing homemade food to COVID-affected families. More than 20k meals, never missed a single day. Didn’t take a single penny from anyone. Thank you to everyone for being a part of this. Stay safe.
Yoga Yoga classes are being conducted for #COVID19 positive people at Sri Padmavathi Nilayam #COVID care center, Tirupati. It helps fight against the mental health challenges of people undergoing treatment. #APFightsCorona #COVID19Pandemic
Donation Hindu temples are doing a great job in these challenging times. Shri Gajanan Maharaj Sansthan, Shegaon, arranged a quarantine facility with 500 beds and daily food for 2,000 people. Money donated and food provided every day!
Blood Donation Kali Puja Blood Donation Camp organized by Ahoban Club - Bangur, Bidhannagar, North 24 Parganas. A noble initiative by Sh. Sajal Bose & the Puja Committee. More than 100 donors donated blood, used for COVID patients.
Table 1: Keywords and Their Association with Negative (Hinduphobic) and Positive & Neutral (Pro-Hindu) Sentiments in Comments
Before Preprocessing After Preprocessing
Rajdeep Sardesai used swastika in his article to portray fascism and nazism to defame Hindus, and Hindus will happily stay silent over such Hinduphobia!! #Hinduphobia #Swastika #Nazism Rajdeep Sardesai used swastika in his article to portray fascism and nazism to defame hindus and hindus will happily stay silent over such hinduphobia
Hindus are the real reason COVID is spreading! Can’t believe they still gather at temples! #HinduSuperspreaders #COVID19 #Pandemic hindus are the real reason covid is spreading cannot believe they still gather at temples hindusuperspreaders covid19 pandemic
Just saw a post about how Hindu rituals are putting everyone at risk! So irresponsible! @username, what do you think? #HinduRituals #COVID19 just saw a post about how hindu rituals are putting everyone at risk so irresponsible [user mention] what do you think hindurituas covid19
LOL, the idea that cow urine cures COVID is just plain ridiculous! #Hinduphobia #Superstitions #COVIDCure laugh out loud the idea that cow urine cures covid is just plain ridiculous hinduphobia superstitions covidcure
OMG, why are they still holding the Kumbh Mela during a pandemic? It’s dangerous and selfish! #KumbhMela #COVID #Pandemic oh my god why are they still holding the kumbh mela during a pandemic it is dangerous and selfish kumbhmela covid pandemic
I suspect Congi ecosystem + rabid Hinduphobia of US news media at play. All US papers (WP, NYT) to Right are unanimous in blaming ”militant H” for this JNU fracas. #USMedia #Hinduphobia i suspect congi ecosystem rabid hinduphobia of us news media at play all us papers wp nyt to right are unanimous in blaming militant h for this jnu fracas usmedia hinduphobia
The thinly veiled Hinduphobia peeping from behind ‘secularism’ facade is tiring and oh so cliche. #Secularism #Hinduphobia the thinly veiled hinduphobia peeping from behind secularism facade is tiring and oh so cliche secularism hinduphobia
Table 2: Examples of tweets related to Hinduphobic before and after preprocessing.
S.No Original Preprocessed S.No Original Preprocessed S.No Original Preprocessed
1 :) happy 7 i’ll’ve I will have 13 @username [user mention]
2 :( sad 8 omg oh my god 14 plz please
3 ain’t am not 9 #COVID19 COVID-19 15 u you
4 lol laugh out loud 10 u2 you too 16 thx thanks
5 smile smile 11 rt retweet 17 idk I don’t know
6 #StaySafe StaySafe 12 asap as soon as possible 18 ik I know
Table 3: Transforming tweet with emojis to standardised language for data processing by language models.
Figure 2: BERT (HP-BERT) Model Architecture for Detecting Hinduphobic Sentiment, Featuring Tokenization (Encoding) of Hinglish/English Tweets Using a Set of Encoders. It Includes a Fully Connected Layer for Classifying Tweets as Hinduphobic or Non-Hinduphobic/Neutral, and the Final Layer Provides Sentiment Classification into Different Categories.
After collecting and curating the Hinduphobic X (Twitter) dataset, we preprocessed it to enhance its quality and relevance by removing irrelevant content, spam, and duplicate entries, along with eliminating web links, hashtags, and profile tags [97, 98, 99]. We also expanded contractions and abbreviations to ensure consistency in data preparation. Additionally, we transformed the emojis into text along with other non-text symbols [100, 101] as shown in (Table 3). Furthermore, (Table 2) presents examples of tweets before and after preprocessing.
We employed a Human-in-the-Loop (HITL) approach [102, 95] to ensure accurate labelling of the processed tweets. We manually labelled 2,000 tweets as either (1) negative sentiment (Hinduphobic) or (2) positive/neutral sentiment (pro-Hindu). These labelled examples were used as training data for the GPT-3.5-Turbo model, accessed via the OpenAI API [^7] https://platform.openai.com/docs/models/gpt-3-5-turbo. After GPT-3.5-Turbo labelled the remaining tweets, the automatically labelled tweets were manually checked. The final labelled dataset, containing 8,000 tweets (4,600 labelled as Hinduphobic and 3,400 as positive/neutral), was published on Kaggle [103]. The need for manual labelling stemmed from the challenge of accurately identifying hate speech, which is language and culture-dependent. Existing models, such as HateBERT [65] and COVID-HateBERT [104], may struggle to detect nuanced expressions of Hinduphobic content with hidden meanings. Correctly identifying instances of hate speech is essential to maintaining data integrity, ensuring accurate interpretation, and minimising the risk of inaccuracies. Figure 1 illustrates the process of dataset creation and processing. This dataset will be used to fine-tune the BERT (HP-BERT) model, a specialised model designed to improve the detection of Hinduphobic content.
3.1.2 SenWave dataset
The SenWave dataset [79], contains over 104 million tweets and Weibo messages related to COVID-19 in six different languages: English, Spanish, French, Arabic, Italian, and Chinese. The data was collected between March 1 and May 15, 2020, to analyse social media posts surrounding the pandemic. It includes 10,000 English and 10,000 Arabic tweets labelled into ten sentiment categories: optimistic, thankful, empathetic, pessimistic, anxious, sad, annoyed, denial, official report, and joking. This comprehensive labelling facilitates in-depth analysis of the diverse emotional responses exhibited during the pandemic. The effectiveness of the SenWave dataset has been validated in two prior studies [71, 105], demonstrating its efficacy for sentiment analysis tasks. After fine-tuning our BERT model (HP-BERT) with the Hinduphobia dataset for Hinduphobic content classification, we will apply multi-stage fine-tuning using the SenWave dataset. Fine-tuning will enable the model ability to perform sentiment analysis on the Global COVID-19 X (Twitter) dataset, refining its understanding of sentiment in the context of Hinduphobic content during the pandemic.
3.1.3 Global COVID-19 X (Twitter) Dataset
The Global COVID-19 X (Twitter) dataset [80] is a comprehensive collection of tweets related to the COVID-19 pandemic, spanning from March 2020 to February 2022. This dataset includes tweets from six countries: Australia, Brazil, India, Indonesia, Japan, and the United Kingdom, as outlined in (Table 4). It captures various public opinions and sentiments about the pandemic, enabling detailed analysis across diverse cultural and regional contexts. This dataset has also been utilised by Chen et al. [17] for sentiment analysis related to Sinophobia (anti-China sentiments) and Chandra et al. [94] for the analysis of antivaccine-related tweets during COVID-19 across different countries. We processed this dataset similarly to the Hinduphobic dataset, ensuring standardised text input essential for the model to accurately capture the contextual nuances in the data. The processed dataset will be used to analyze Hinduphobic tweets across different countries, as well as sentiment analysis in various regions using our HP-BERT model.
Country Name Number of Tweets
Australia 3,212,464
Brazil 943,913
India 5,411,294
Indonesia 229,935
Japan 644,510
United Kingdom 16,958,866
Table 4: Tweet counts from six different countries.
Figure 3: A multi-stage process for HP-BERT model training and analysis: Stage 1 involves dataset collection, Stage 2 preprocesses data, Stage 3 fine-tunes HP-BERT for binary and sentiment classification, Stage 4 applies classification and trend analysis, and Stage 5 conducts sentiment analysis with visualisation of trends and polarity.
3.2 Hinduphobic-BERT Model
BERT [55] is an advanced language model designed to understand the contextual relationships between words, by simultaneously considering both left and right contexts in a sentence. BERT uses a bidirectional attention mechanism based on the Transformer architecture [106], allowing it to read an entire sentence at once and understand the context more effectively. The attention mechanism enables BERT to weigh the importance of different input parts, making it superior for language understanding tasks such as sentiment analysis [107]. The role of BERT in this context is to generate contextualised embeddings for each tweet in the dataset. The pre-training on a massive corpus enables BERT to capture nuanced language features, making it highly suitable for sentiment analysis tasks [94], and detecting hate speech [108] and bias patterns. Furthermore, the BERT model can be easily extended to domain-specific tasks by fine-tuning with specialised datasets.
We develop a Hinduphobic BERT (HP-BERT) model to classify tweets containing Hinduphobic sentiments. The HP-BERT model architecture for detecting Hinduphobic sentiment is shown in (Figure 2). The process begins with the Hinduphobic BERT Tokenizer, which tokenises Hinglish (a blend of Hindi and English), and English tweets into input identifiers (IDs), input type IDs, and input mask. Input IDs represent the tokenised text in numeric form, enabling the model to process it, while input type IDs distinguish between different parts of the input, such as multiple sentences. The attention mask ensures that the model focuses only on meaningful tokens, ignoring padding. These inputs are then passed through a 12-layer encoder to generate contextualised embeddings that are later fed into the fully connected to classify each tweet as either Hinduphobic or Non-Hinduphobic/Neutral. We then use an additional layer for sentiment classification by training (refining) using the SenWave dataset [79]. By combining Hinduphobic detection with sentiment classification, the HP-BERT model enables a comprehensive sentiment analysis of tweets, which also includes the computation of a sentiment polarity score based on predefined weights in (Table 5). This approach allows the model to detect Hinduphobic content while providing detailed sentiment categorisation and sentiment polarity analysis.
Sentiment Weight
Optimistic 2
Thankful 3
Empathetic 0
Pessimistic -4
Anxious -2
Sad -3
Annoyed -1
Denial -5
Official 0
Joking 1
Table 5: Weight ratios for different sentiments for polarity score calculation.
Data Method Accuracy (Mean $\pm$ Std) F1 (Mean $\pm$ Std) Precision (Mean $\pm$ Std) Recall (Mean $\pm$ Std)
Hinduphobic Dataset (20/80) BERT 0.9053 $\pm$ 0.0051 0.9048 $\pm$ 0.0053 0.9059 $\pm$ 0.0050 0.9050 $\pm$ 0.0054
Hinduphobic Dataset (40/60) BERT 0.9215 $\pm$ 0.0042 0.9210 $\pm$ 0.0044 0.9221 $\pm$ 0.0041 0.9213 $\pm$ 0.0045
Hinduphobic Dataset (60/40) BERT 0.9426 $\pm$ 0.0035 0.9421 $\pm$ 0.0037 0.9430 $\pm$ 0.0034 0.9423 $\pm$ 0.0038
Hinduphobic Dataset (80/20) BERT 0.9472 $\pm$ 0.0028 0.9465 $\pm$ 0.0030 0.9474 $\pm$ 0.0029 0.9467 $\pm$ 0.0031
Table 6: Evaluation metrics (mean and standard deviation) for HP-BERT across 30 experimental runs. Metrics include accuracy, F1 score, precision, and recall for various training/testing splits on Hinduphobic COVID-19 X (Twitter) Dataset.
4 Framework
Our Hinduphobic tweet detection and sentiment analysis framework involves multiple stages and components, as illustrated in (Figure 3).
In Stage 1, we obtain three datasets to fine-tune and evaluate our HP-BERT model for sentiment classification, including the i. Hinduphobic COVID-19 X (Twitter) Dataset [103], ii. SenWave Dataset [79], and iii. Global COVID-19 X (Twitter) Dataset [80]. In the case of Hinduphobic COVID-19 X (Twitter) Dataset, we use HITL for manual labelling of 8000 tweets and also develop a model for classification of the leftover tweets.
In Stage 2, we preprocess the three datasets by removing irrelevant content, spam, and duplicate entries, as well as eliminating web links, hashtags, and profile tags. We also expanded contractions and abbreviations to ensure consistency across the data. Additionally, we convert the emojis to text along with other symbols for expressions used in social media.
In Stage 3, we first fine-tune a pre-trained BERT model using the ”Hinduphobic COVID-19 X (Twitter) Dataset” to detect and classify Hinduphobic content in tweets. Afterwards, we further fine-tune the model on the ”SenWave dataset” to incorporate sentiment classification with an additional layer, as shown in (Figure 2). Another layer is added for sentiment classification, allowing the model to extract sentiments associated with the tweet to enable the HP-BERT model to perform both Hinduphobic content detection and sentiment analysis (multi-label sentiment classification).
In Stage 4, we utilise the trained HP-BERT model for ”the Global COVID-19 X (Twitter) Dataset” to examine the volume of Hinduphobic content by examining various dimensions: total counts, country-wise distribution, and month-wise trends. Additionally, we conduct bigraph, bigram, and trigram analyses both globally and individually for each country. This multi-faceted analysis helps us understand how Hinduphobic content evolved and varied geographically. The insights gained from these analyses provide a valuable understanding of global patterns and trends in Hinduphobic discourse during the pandemic. Additionally, we apply HateBERT [65] to assess its effectiveness in detecting hate speech and abusive content when related to Hinduphobia.
In Stage 5, we apply our HP-BERT model to perform sentiment analysis on Hinduphobic content, which involves conducting a longitudinal analysis and generating various plots to examine the sentiment distribution and trends across different regions, including sentiment polarity scores. Polarity scores are calculated using the TextBlob library, and custom weight ratios for different sentiment labels as shown in (Table 5) are applied to enhance the analysis. This stage provides deeper insights into how sentiments in Hinduphobic content fluctuate over time and vary geographically.
5 Results
5.1 Technical setup
We use a dropout regularisation rate of $d=0.2$ to prevent overfitting in BERT, obtained from Huggingface [^8] https://huggingface.co/docs/transformers/en/model_doc/bert. The model is trained for 10 epochs with a learning rate of $2\times 10^{-5}$ and a batch size of 8. Additionally, weight decay regularisation with a rate of $\gamma=0.01$ is applied to further mitigate overfitting. We utilise Python-based libraries, including Transformers, NumPy, Pandas, Matplotlib, and PyTorch. The training process is accelerated using an NVIDIA GeForce RTX 2080 Ti GPU (graphics processing unit)with 11264 MiB of memory and an AMD Ryzen 7 5800H processor, along with 16 GB of system memory.
5.1.1 Fine-tuning the HP-BERT Model
We fine-tuned the HP-BERT model for Hinduphobia detection and sentiment analysis using the ”Hinduphobic COVID-19 X (Twitter) Dataset”, which contains 8,000 labelled tweets. The dataset was split into four different training/testing configurations (randomly). The fine-tuned model was evaluated on the test set to assess its performance in classifying Hinduphobic tweets. (Table 6) presents the results from 30 independent experimental runs, reporting the mean and standard deviation for accuracy, F1 score, precision, and recall. The results demonstrate strong performance, with high mean accuracy and balanced F1 scores, indicating the model’s ability to maintain precision and recall. Next, we used the fine-tuned HP-BERT model to predict Hinduphobic tweets in a global dataset.
In (Table 7), we present the results of the fine-tuned HP-BERT model on the SenWave dataset for sentiment analysis, offering additional analysis following Hinduphobic tweet detection. Since this is a multi-label sentiment analysis task, conventional metrics are not applicable. Instead, we report the Label Ranking Average Precision (LRAP) score, F1-score, Jaccard score, and Hamming loss. These metrics have been widely used in earlier studies [94, 71], and our results further validate their findings.
Metric Value
Hamming Loss 0.1476
Jaccard Score 0.5013
Label Ranking Average Precision (LRAP) 0.7501
F1 Score (Macro) 0.5187
F1 Score (Micro) 0.5834
Table 7: Evaluation metrics of the HP-BERT model after fine-tuning on the SenWave dataset for sentiment analysis.
5.2 Data Analysis on Global COVID-19 X (Twitter) Dataset)
5.2.1 Analysis of COVID-19 Cases
We first review the COVID-19 case trends from April 2020 to January 2022 across six countries, as shown in (Figure 4). We notice that India had a major peak in cases in April-May 2021, likely due to the Delta variant, while Brazil and Indonesia also saw significant peaks in mid-2021. The United Kingdom shows a rise toward late 2021, likely from the Omicron variant. Japan and Australia had relatively low case numbers throughout, with slight increases at the end of 2021. Overall, the trends reflect the varied impact of COVID-19 waves, and variants spread across these regions.
Figure 4: Monthly trends in COVID-19 case counts.
Figure 5: Monthly trends of Hinduphobic tweets across all countries.
Figure 6: Monthly trends of Hinduphobic tweets across countries excluding India.
Dataset Method Australia Brazil India Indonesia Japan UK
The Global Dataset HateBERT [65] 137699 36734 151659 8109 26518 64121
The Global Dataset HP-BERT 479 352 9242 356 389 746
Subset of The Global Dataset HateBERT [65] 48 33 737 39 29 65
Table 8: Tweet counts across six countries using HateBERT and HP-BERT on the global dataset and its subset.
5.2.2 Results of HateBERT model
We applied HateBERT [65] to analyze hate speech and abusive content during the COVID-19 pandemic across six countries, with the results summarised in (Table 8). To gain deeper insights into recurring themes within the discourse, we performed bigram and trigram analyses, as shown in (Figure 7 Panel-(a)) shows the most frequent bigrams, which reveal key patterns in the pandemic-related discourse. Bigrams such as ”free vaccine” and ”pvt hospital” highlight concerns around vaccine accessibility and healthcare disparities, while ”modi govt” and ”crore job” reflect criticisms of the Indian government and the economic challenges faced by citizens. (Figure 7 Panel-(b)) presents the most frequent trigrams, offering further insights into more complex interactions within the conversations. Trigrams such as ”daily wage labourer” and ”crore job loss” underscore the economic vulnerabilities faced by workers during the pandemic, while ”lockdown announced train” and ”stopped interstate travel” emphasise the disruptions caused by lockdown measures, particularly on mobility and commerce.
Although the results from HateBERT provide valuable general insights into pandemic-related discourse, they remain broad. To further refine our analysis and specifically focus on Hinduphobic content, we applied the HP-BERT model in subsequent analyses.
(a) Bigram
(b) Trigram
Figure 7: Bigram and trigram analysis of tweets labelled by HateBERT worldwide.
5.2.3 Results HP-BERT model
Using the HP-BERT model, we analyse Hindubhobic tweets and associated sentiments across countries and regions, providing insight into the perceptions of Hindu communities in various parts of the world during and after COVID-19. This can enable us to identify sentiment trends and potentially inform strategies for addressing harmful stereotypes and communal biases on social media platforms.
Figure 8: Monthly distribution of Hinduphobic tweets identified by the HP-BERT model.
Finally, we applied our HP-BERT model to the Global COVID-19 X (Twitter) Dataset to analyze Hinduphobic tweets. (Figure 5) presents a multi-graph chart illustrating tweet counts across different countries, showing the monthly distribution of Hinduphobic tweet counts from April 2020 to January 2022. (Table 8) provide a detailed breakdown of the tweet counts across six countries where India recorded the highest number of Hinduphobic tweets, accounting for over 9,242 tweets, significantly surpassing other countries such as the United Kingdom, Australia, Japan, Indonesia, and Brazil, which reported much lower tweet counts. This higher number in India can be attributed to its large Hindu population [109], along with the presence of other significant religious communities such as Muslims and Christians. The United Kingdom ranked second, followed by Australia, Japan, Indonesia, and Brazil. This analysis highlights that Hinduphobic sentiment on Twitter during this period was predominantly concentrated in India. According to the 2011 Indian census [110], Hinduism encompassed 79.8% of the population followed by 14.2% Islam, and 2.3% Christianity.
We created an additional visualisation excluding India, as shown in Figure 6 to allow for a clearer view of Hinduphobic tweet counts across the remaining countries. We also present a distribution of Hinduphobic tweets (Figure 8) to review the fluctuating levels of Hinduphobic tweets, with notable peaks (April to May 2020 and April to May 2021). After these peaks, we can observe a general downward trend with periodic increases in March to April 2021, and September 2021. This distribution suggests varying intensities of Hinduphobic sentiment on Twitter throughout the observed period, with some months experiencing sharp rises in tweet volume while others show a noticeable decline.
(a) Bigram
(b) Trigram
Figure 9: Bigrams and trigrams extracted from Hinduphobic tweets identified by HP-BERT worldwide.
We performed n-gram analysis [111] to explore the tweets processed by HP-BERT, focusing on bigrams and trigrams. (Figure 9) illustrates the top 10 bigrams and trigrams extracted from tweets posted between April 2020 and January 2022 across six countries. In (Figure 9 Panel-(a)), the most frequent bigrams in the dataset include ”covid 19”, ”cow urine” and ”cow dung”. Notably, the bigram ”cure covid” appears frequently, suggesting widespread discussions mocking Hindus for purportedly using cow urine and cow dung as COVID-19 treatment due to the sacred nature of cows in Hinduism. Other bigrams, such as ”anti hindu”, ”hindu temple” and ”spread covid” highlight tweets blaming Hindus for spreading COVID-19 by attending temples. Furthermore, phrases such as ”hinduphobia hindu”, ”hinduphobia hinduphobia” and ”hinduphobia real” contribute to the growing narrative of negative sentiment against Hindus.
(Figure 9 Panel-(b)) presents the most frequent trigrams, including ”cow urine cure”, ”drink cow urine”, ”urine cure covid”, ”cow dung urine” and ”cure covid 19”. These terms reflect stereotypes about Hindus using cow-related products to combat COVID-19, as seen in reports of events such as cow urine-drinking gatherings to ward off the virus [112] and such narratives have contributed to discrimination against Hindus in several countries [113]. Additionally, trigrams like ”dgh conf agenda” and ”conf agenda hinduphobia” refer to a conference held in the United States aimed at addressing global Hinduphobia [^9] https://shorturl.at/8xCUW. Phrases such as ”spread covid 19” and ”dismantling global hindutva” explicitly link Hindus to the spread of COVID-19, further fuelling anti-Hindu sentiments worldwide.
(a) Australia
(b) Brazil
(c) India
(d) Indonesia
(e) Japan
(f) United Kingdom
Figure 10: Top 10 bigrams and trigrams of Hinduphobic tweets identified by HP-BERT for different countries.
We also extracted the top 10 bigrams and trigrams for each of the six countries, as shown in (Figure 10). We find that certain bigrams, such as ”covid 19”, ”cow urine” and ”cow dung” are consistently present across all six countries. Similarly, the most frequent trigrams are common across countries, aligning with global patterns. These recurring phrases reflect a shared narrative that transcends national boundaries, emphasising the widespread dissemination of stereotypes and sentiments linked to Hinduphobia.
In Australia, notable bigrams include ”hindu temple”, ”hinduphobia hindu”, ”anti hindu”, ”folded hand” (a reference to the Indian greeting ”Namaste”), and ”abrahamic faith” often used to mock Hindu practices and promote a sense of religious superiority. Trigrams such as ”hindu temple attacked”, ”anti hindu bigotry” and ”abrahamic faith supremacy” highlight themes of discrimination and religious intolerance. In Brazil, key bigrams include ”cure covid”, ”hinduphobia problem”, ”abrahamic faith”, and ”hindu temple”. The trigrams ”like bbc hinduphobia”, ”bbc hinduphobia problem”, ”urine cure covid”, ”drink cow urine” and ”cure covid 19” indicate mockery directed toward Hindu practices, particularly regarding traditional remedies. In India, prominent bigrams include ”hinduphobia real”, ”anti hindu”, and ”hinduphobia hindu”. Trigrams such as ”drink cow urine”, ”cow urine cure”, ”cure covid 19”, and ”dismantling global hindutva” underscore narratives aimed at ridiculing Hindu practices and beliefs. For Indonesia, bigrams such as ”hindu temple”, ”hinduphobia hindu”, ”skin tone” and ”folded hand” point to cultural and racial stereotyping. The trigrams such as ”hindu temple attacked”, ”abrahamic faith supremacy”, ”cow dung urine” and ”religion called intolerant” suggest mockery of Hindu traditions and communal narratives. In Japan, bigrams such as ”cure covid”, ”stop hinduphobia”, and ”attack hindu” reflect growing concerns about Hinduphobia. The trigrams such as ”cure covid 19”, ”drink cow urine”, ”hindu temple attacked” and ”spread covid 19” emphasise the global spread of derogatory narratives targeting Hindus. Lastly, in the United Kingdom, bigrams like ”hindu temple”, ”anti hindu” and ”hinduphobia real” are significant. Trigrams such as ”drinking cow urine”, ”hindu temple attacked”, ”spread covid 19” and ”cow dung urine” further illustrate how stereotypes are perpetuated. Overall, while common themes of Hinduphobia are observed across countries, specific regional narratives and stereotypes are evident in the analysis. These findings reveal how Hindu practices, beliefs, and culture are frequently mocked or misrepresented, reflecting shared global patterns and localised biases.
5.3 Analysis of HP-BERT results using HateBERT
Next, we apply HateBERT [65] to the results of HP-BERT to assess its effectiveness in detecting hate speech and abusive content, particularly in the context of Hinduphobia across six countries (Table 8). The detection results were comparatively lower because HateBERT is primarily trained on English language data, which may not effectively capture nuanced and culturally specific Hinduphobic rhetoric. HateBERT was pre-trained on a corpus of Reddit comments, making it more effective at detecting overtly hateful and abusive content. Furthermore, we generate bigrams and trigrams as shown in (Figure 11) to highlight recurring themes. These phrases reflect recurring themes that highlight narratives of Hinduphobia, particularly in the context of COVID-19 remedies, cow-related references, and religious or cultural stigmas. This analysis underscores the prevalence of stereotypes and how they manifest across different countries.
(a) Bigram
(b) Trigram
Figure 11: Bigrams and trigrams of Hinduphobic tweets identified by HP-BERT analyzed using HateBERT worldwide.
5.4 HP-BERT Sentiment Analysis
We next employ the HP-BERT to perform sentiment analysis on Hinduphobic tweets, capturing a range of sentiment categories. (Table 7) provides evaluation metrics such as accuracy, precision, recall, and F1-score. Our analysis reveals a wide range of sentiments expressed in the tweets. We also examined user networks, activity patterns, and profile descriptions to identify common traits among users sharing Hinduphobic content. These insights offer a deeper understanding of the context and spread of Hinduphobia across social media platforms.
Figure 12: Percentage of Hinduphobic tweets with a different number of labels.
Figure 13: Number of Hinduphobic tweets with different numbers of labels
Figure 14: Total number of Hinduphobic tweets for each sentiment.
(Figure 12) illustrates the distribution of sentiment labels assigned to each tweet in the dataset of Hinduphobic tweets from various countries, including Australia, Brazil, India, Indonesia, Japan, and the United Kingdom. The pie chart indicates that 0.2% of the tweets have no sentiment label, 80.9% have one sentiment label, 15.8% have two sentiment labels, and 3.1% have three or more sentiment labels. (Figure 13) provides a detailed view of the number of sentiment labels assigned to tweets across different countries. The bar graph reveals a highly skewed distribution of tweets by the number of labels across all the countries. India dominates with the largest number of tweets, particularly in the one-label category, while other countries, including Australia, Brazil, Indonesia, Japan, and the United Kingdom, have significantly lower tweet counts across all label categories. Tweets with a single label are most common, while those with multiple labels (one or two sentiments) are much fewer, highlighting the concentration of data in simpler classifications and indicating a more complex sentiment expression in these regions.
Figure 15: Heatmap showing the distribution of Hinduphobic tweets.
We next conduct a longitudinal analysis of sentiments expressed in tweets across different countries from April 2020 to January 2022. Our initial focus was on the total number of tweets associated with each sentiment, as illustrated in (Figure 14). The sentiment ”annoyed” emerged as the most prevalent, with over 9,000 tweets followed by ”official report” which has a much lower count. The ”joking” sentiment is also prominent, particularly in tweets mocking Hindus. We also find ”optimistic”, ”denial”, ”anxious” and ”sad” and in contrast, sentiments such as ”thankful”, ”pessimistic” and ”empathetic” appeared in a very small number of tweets, with occurrences nearing zero. The heatmaps in (Figure 15) illustrate the frequency of co-occurring sentiments in the extracted tweets for the years 2020 and 2021. The most frequently observed individual sentiment was ”annoyed” followed by ”joking” and ”denial.” The highest co-occurrence was noted between ”annoyed” and ”joking”, indicating that these sentiments are often expressed together.
Figure 16: Percentage of tweets of different sentiments for each country.
We next exclude the ”official report” category from the rest of the analysis to focus on the emotional content of the tweets for a country-wise examination to gain deeper insights into the sentiment distribution (Figure 16 Panel-(a)). The results show that ”annoyed” is the most dominant sentiment, with the highest percentage across all countries. The ”joking” sentiment consistently ranks second across all countries but appears at a significantly lower percentage compared to ”annoyed”. The ”optimistic” sentiment is also present at a lower level, however, in Japan, we observe nearly equal to ”annoyed”. The ”thankful” sentiment is scarcely present, with percentages approaching zero in all countries. (Figure 16 Panel-(b)) presents the distribution of ”empathetic”, ”pessimistic”, ”sad”, ”anxious” and ”denial” sentiments; where ”denial” emerges as the most frequently occurring sentiment across all countries. Other negative sentiments appear in minimal quantities across all countries. In summary, ”denial” reflects the prevailing negative sentiment against Hindus in the six countries, despite its relatively low frequency.
5.5 Polarity Score
We employ TextBlob and a custom weight strategy for polarity scores to quantify the sentiment expressed in text, categorising it as positive, negative, and neutral, where scores close to zero represent neutral sentiment. TextBlob [114] is a widely used Python library that provides various tools for NLP tasks, including sentiment analysis. The custom weight strategy involves assigning specific weights to each sentiment category (Table 5) to capture more complex sentiment expressions that traditional sentiment analysis tools may overlook. Figure-17 Panel-(a) presents the TextBlob polarity score revealing that a substantial portion of tweets has polarity scores close to zero, indicating a predominantly neutral sentiment. Furthermore, Figure 17 Panel-(b) presents the custom weight strategy, which demonstrates a concentration of scores near zero. However, the custom approach provides a more detailed representation, revealing multiple distinct peaks at various negative scores, notably at -0.20, which suggests more effectiveness in capturing subtle variations in sentiment.
(a) TextBlob
(b) Custom weights
Figure 17: Distribution of TextBlob and Custom weight polarity scores.
(a) TextBlob polarity
(b) Custom weight polarity
Figure 18: Mean polarity of TextBlob and custom weight over month.
(Figure 18) presents the mean polarity scores calculated monthly using TextBlob and a custom-weighted strategy for six countries spanning April 2020 to January 2022. In (Figure 18 Panel-(a)), TextBlob polarity scores show trends in sentiment over time, with a notable positive spike around May 2020, coinciding with early COVID-19 response measures that may have fostered optimism and solidarity. As the pandemic progressed, the sentiment generally became more negative, reflecting challenges such as pandemic fatigue, economic uncertainty, and health crises. Trends varied by country: Australia exhibited pronounced fluctuations, while India and Indonesia showed relatively stable, neutral sentiments. Brazil and Japan experienced notable swings, likely influenced by differing public reactions or media narratives. The custom-weighted polarity scores, shown in (Figure 18 Panel-(b)), display greater sensitivity to negative sentiment compared to TextBlob. A significant dip in polarity scores is observed around May 2021, coinciding with a surge in COVID-19 cases, particularly in India. This indicates heightened negativity associated with the pandemic’s severe impact. Conversely, countries with relatively stable case numbers, such as Japan and Australia, exhibited fewer fluctuations in sentiment scores. The rise in polarity scores in November 2021 may be linked to shifting public attention from the origins of COVID-19 to vaccine distribution and recovery efforts. As discussions shifted towards vaccination progress and pandemic management, the overall sentiments changed globally, reflecting a more optimistic outlook.
6 Discussion
Our study analyses Hinduphobic tweets and their sentiment dynamics during the COVID-19 pandemic. We gained valuable insights into how public sentiment evolved during this period and how Hinduphobia manifested in different contexts for the selected countries. We first analysed the trends in COVID-19 cases (Figure 4) and their correlation with the prevalence of Hinduphobic tweets. As shown in (Figures 8) monthly distribution of Hinduphobic tweets and (Figure 5) country-wise distribution, peaks in COVID-19 cases often coincided with spikes in Hinduphobic tweets. This trend is further reflected in (Figure 18), which shows a drop in mean polarity scores, particularly during major outbreaks and significant public health interventions. These findings suggest a strong link between pandemic-related stressors and the rise in negative sentiment and Hinduphobic discourse.
Our HP-BERT model predictions for Hinduphobic tweets, as presented in (Table 8), show the highest number of Hinduphobic tweets originating from India, followed by the United Kingdom, Australia, Japan and Brazil. For sentiment analysis (Figure 14), the emotions ”annoyed” and ”official report” were common, driven by tweets sharing anti-Hindu narratives, such as criticism of traditional practices, targeting Hindu festivals, or reacting to perceived bias in official reports. Our analysis highlighted a predominant presence of negative sentiments towards Hindus, with ”annoyed” and ”official report” being the most common emotions. These findings align with previous studies, such as Wang et al. [17] on Sinophobia sentiment analysis, and Chandra et al. [71, 105] on pandemic-related sentiment analysis, underscoring the robustness of sentiment analysis models in effectively capturing predominant emotions across diverse datasets. The analysis of the bigrams and trigrams in (Figure 9) for all countries and (Figure 10) for country-wise trends, revealed patterns in Hinduphobic tweets. These patterns are driven not only by the stereotypical association of Hindus with the coronavirus but also by a complex interplay of health, racial, and political factors. The frequent use of politically charged terms underscores how the discourse surrounding COVID-19 and its origins became highly politicised, with political narratives significantly influencing public sentiment during the pandemic.
We also analysed the results of the HP-BERT model, as shown in (Table-8), and by using HateBERT [65] to assess its effectiveness in detecting hate speech and abusive content, particularly in the context of Hinduphobia across six countries. The detection results were lower for HateBERT because it is primarily trained on English-language Reddit data, which may not capture the nuanced, culturally specific Hinduphobic rhetoric. While HateBERT is more effective at identifying overtly hateful content, it struggles with the subtleties of Hinduphobia. Additionally, the bigrams and trigrams shown in (Figure 11) highlight recurring themes of Hinduphobia, particularly related to COVID-19 remedies, cow-related references, and religious stigmas, emphasising the persistence of stereotypes across countries. The analysis of sample tweets, as shown in Table 9 and (Table 10), indicates that the model performs well in accurately classifying sentiments in most cases. The samples generally align with expected sentiment patterns, showcasing the model’s effectiveness in capturing the predominant sentiments expressed in the tweets. However, specific examples reveal challenges in sentiment analysis, particularly those highlighted by Hussein [115], such as handling negation, sarcasm, and context-dependent language. For instance, consider the second tweet in (Table 9) ”Hindu Temples have donated huge sums of money so that people get food and all this is good”. This tweet was classified as ”Optimistic” with a positive polarity score of ”0.2,”. Similarly, in (Table 10), the tweet ”RT @NewsHandle: Religious events like Kumbh is reckless and spreading the virus. Why does India allow this?” was labelled as ”critical” and ”official report” with a polarity score of ”-0.3.” Although the classification of ”critical” aligns with the sentiment expressed, it fails to capture the underlying bias selectively targeting Hindu practices, highlighting the need for more context-sensitive sentiment analysis. These examples underscore the complexities of interpreting tweets with subtle or implicit biases. In the context of Hinduphobia, sarcastic, ironic, or biased language can easily evade detection or be misclassified. This highlights the importance of refining sentiment analysis models to address these challenges and ensure a more accurate understanding of the nuanced expressions of Hinduphobia in social media discourse.
Hinduism is a religion that encompasses diverse communities both in ethnicity and belief system with a pluralistic and inclusive tradition that combines pantheism, polytheism, and non-theism [116, 117]. The core texts encompass the Vedic literature such as Bhagavad Gita that lay the foundation of Sanatan Dharma, which teaches respect and acceptance for all beliefs and cultures without a requirement of religious conversion. This philosophy is embodied in the ancient concept of Vasudhaiva Kutumbakam, meaning (the whole world is one family) [118, 119]. Rooted in compassion, non-violence, and universal unity, Sanatan Dharma encourages a harmonious existence where every being is interconnected, celebrating diversity and embracing humanity as one vast family [120].
The dark historical practices associated with Hinduism, such as the caste system, sati (widow burning), child marriage, dowry, untouchability, temple entry restrictions, gender-based restrictions, animal sacrifice, the devadasi tradition, and caste hierarchy have long been reformed, abolished, and rarely practised in modern Hindu communities. The caste system [121], for instance, originated as a division of labour, but evolved into a rigid social hierarchy; however, it has been legally abolished by India’s Constitution and affirmative actions were introduced to promote equality. The practices such as sati [122] and child marriage [123] were outlawed in the 19th and early 20th centuries, and the dowry system [124] was criminalised in 1961. Temple entry [125] and gender-based restrictions [126] have largely been lifted due to reform movements, while animal sacrifice [127] and the devadasi system [128] have also seen significant cultural and legal rejection. Despite these advancements, Hindu society has often been misrepresented in Western media, which sometimes portrays these now-abolished practices as if they were active elements of Hinduism today [129, 130]. Western media still talks about outlawed practices as if they are part of Hinduism today [131, 132], without understanding the progress that has been made. This leads to a misleading view of Hinduism, showing it as stuck in the past, while ignoring the positive changes and efforts made to promote equality and social justice. By focusing on these old issues [133], the media overlooks the more inclusive and modern aspects of Hinduism, creating a negative image based on misunderstandings.
In ancient times, Hinduism was widespread and in the majority across many regions, but over centuries, political changes, invasions, and the spread of other religions led to a significant decline in the Hindu population in several countries [134, 135, 136]. In Pakistan [137, 138] Hinduism was once the dominant religion, especially in regions like Sindh and Punjab, with the Indus ValleyCivilisationn also having Hindu-related practices. However, where Hindus likely made up ($60\%-70\%$) of the population. Today, Hindus make up only ($1.85\%$) of Pakistan’s population [^10] https://en.wikipedia.org/wiki/Hinduism_in_Pakistan, which is due to forced conversions and low birth rates. In Bangladesh, Hinduism [139, 140] was once the dominant religion, especially under the Pala dynasty, with Hindus constituting around ($50\%-60\%$) of the population and now, Hindus make up about ($8\%-10\%$) of the population [^11] https://en.wikipedia.org/wiki/Hinduism_in_Bangladesh. In Afghanistan [141, 142], Hinduism was widely practised, especially in the Gandhara region, where Hindus were about ($40\%-50\%$) of the population, but today, only about ($0.01\%$) remain. Indonesia [143, 144] was also once predominantly Hindu, particularly during the Majapahit Empire, with Hindus making up around ($70\%-80\%$) of the population, but now only ($1.7\%$) of the population follows Hinduism, with Bali as the exception. The Maldives [145] was once a Hindu-majority region, with nearly ($90\%-100\%$) Hindus before the 12th century, but today, less than ($0.01\%$) of the population remains Hindu. In Sri Lanka [146, 147] Hinduism was widespread, especially in the Tamil-majority northern regions, where it was around ($30\%-40\%$) of the population in ancient times, but today, Hindus make up about 12% of the population. This shows how Hinduism, once dominant in these areas, has now become a minority religion in most of them.
The decline of the Hindu population in these regions is due to a mix of political, social, and religious factors. In Pakistan and Bangladesh, the spread of Islam, invasions, forceful conversions, and marriages of Hindu girls to Muslim men led to a loss of Hindu identity. Discrimination, social exclusion, and restrictions on Hindu practices also contributed [148]. In Afghanistan, Islam pushed Hinduism to the margins, reducing its population. In Indonesia and the Maldives, Islam’s spread led to the near disappearance of Hinduism, except in Bali. In Sri Lanka, Buddhism became dominant, and Hinduism declined due to forced assimilation and restrictions. In the Middle East, Hindus faced similar discrimination as Islam spread, with forced conversions and the suppression of Hindu rituals, causing a decline in Hindu communities in regions like ancient Persia and the Arabian Peninsula [149].
The lack of education about the scientific and philosophical heritage of Hinduism, both within India and globally, exacerbates the ridicule of its achievements. Some of the prominent Hindu philosophers such as Adi Shankaracharya (788–820 CE), Swami Vivekananda (1863–1902), Ramanuja (1017–1137 CE), and Madhvacharya (1238–1317 CE), along with mathematicians/scientists such as Aryabhata (476–550 CE), Srinivasa Ramanujan (1887–1920), Bhaskaracharya (Bhaskara II) (1114–1185 CE) made monumental contributions to philosophy, mathematics, astronomy, and science. Modern thinkers and educators, including Mahatma Gandhi (1869–1948) and Dr. B. R. Ambedkar (1891–1956) [150, 151] also significantly influenced intellectual and social thought. Despite their profound contributions, they have been underrepresented in educational curricula in India and abroad. This lack of recognition perpetuates a distorted narrative that dismisses ancient Indian knowledge, overshadowed by colonial biases. Internationally, the limited exposure to Hindu philosophy and science further reinforces stereotypes, creating a cultural vacuum filled with ridicule rather than respect.
This study has several limitations that should be acknowledged. First, all three datasets were primarily derived from X (Twitter), which may not fully capture the diversity of public opinion on Hinduphobia. The platform’s user base skews younger, more urban, and technologically savvy tweeters [152] with potential demographic biases. Additionally, the 280-character limit restricts the depth and context of expressed sentiments, making it difficult to capture nuanced perspectives. Wankhade et al. [153] reported that the informal language, abbreviations, and emojis further complicate sentiment analysis, reducing the precision of models such as our HP-BERT and HateBERT. These constraints highlight the need for incorporating additional data sources and methodologies to gain a broader understanding of Hinduphobia. Second, our HP-BERT model struggled with recognising sarcasm, irony, and negation, all critical for accurate sentiment interpretation. Nevertheless, sarcasm is also difficult to be recognised by humans and also has cultural biases that affect its understanding. It also faced challenges in identifying spam and fake content, which could undermine the reliability of classification. Despite efforts to filter and clean the dataset, noisy data such as irrelevant tweets or mislabelled sentiments remains an issue, distorting analysis and reducing accuracy.
Another significant limitation lies in the challenges faced by LLMs in monitoring racist slurs and discriminatory remarks, especially in a multilingual and multi-religious context like India. Social media often combines languages, such as ”Hinglish” (a mix of Hindi and English in Romanized text) [81], further complicating detection. As Agnihotri et al. [154] highlight, allusions, metaphors, and creative language constructs make it difficult to detect implicit personal attacks or Hinduphobic slurs. Although platforms have mechanisms to remove overt hate speech, subtler forms often evade detection. For example, tweets like ”Religious gatherings are the reason the virus is spreading” or ”India’s obsession with ancient cures won’t help in a pandemic” may appear as critiques but often carry implicit biases against Hindu practices. Similarly, ”This second wave is karma for how they treat minorities” politicises the crisis while promoting anti-Hindu narratives. These examples demonstrate the need to refine sentiment analysis models to better handle nuanced and context-dependent language, ensuring a more accurate classification of potentially harmful content.
In future research, the scope of data sources can be expanded beyond Twitter to include other social media platforms such as Facebook, Instagram, Reddit, 4chan (known for extremist and controversial discussions), VK (a Russian platform with less moderation), and platforms such as Parler, Gab, and Weibo. Each of these platforms has distinct user demographics and moderation policies, making their inclusion valuable for capturing a broader spectrum of opinions and emotions. Incorporating data from these sources would provide a more comprehensive understanding of public sentiment, which can be extended to improving models for other problems such as Sinophobia [17]. Additionally, while BERT-based models have proven robust in sentiment analysis, future work could explore the application of more advanced large language models [155, 156]. As the internet is dominated by GenZ users (those born between the mid-1990s and early 2010s), who often use GIFs, emojis, and memes to express their opinions [157], we need to better account for non-textual and visual forms of communication. Humour, irony, and implicit messages in memes represent unique challenges in sentiment analysis that require innovative approaches. Future studies could utilise irony detection models and compare their performance with other sentiment analysis models. This comparison would help identify the strengths and limitations of various approaches, informing the development of more sophisticated tools. Longitudinal studies could also be conducted to analyze sentiments in entertainment media, such as movies and music, over the past few decades to identify trends and shifts in public sentiment. Finally, future sentiment analysis should integrate cultural, religious, historical, political, and economic factors, as these significantly influence public sentiment. For instance, during the COVID-19 pandemic, these factors played a pivotal role in shaping narratives and biases, including the rise of Hinduphobia. Events such as the portrayal of religious gatherings as superspreaders or the misrepresentation of traditional practices contributed to biased narratives. Therefore, future research can provide a more nuanced and comprehensive understanding of the drivers behind Hinduphobic sentiments, offering deeper insights into how such biases evolve and are sustained during global crises.
7 Conclusion
The COVID-19 pandemic has significantly amplified Hinduphobic sentiments, resulting in widespread stereotyping and discrimination against the Hindu community. Our study employed LLMs to perform sentiment analysis on social media data, with a focus on Hinduphobic narratives during the pandemic. We curated a Hinduphobic COVID-19 X (Twitter) Dataset and fine-tuned the Hindupbobic-BERT model to review the evolution of Hinduphobic sentiments over time, including COVID-19. The HP-BERT model demonstrated moderate effectiveness with an additional sentiment analysis component using SenWave dataset.
The analysis revealed a correlation between spikes in Hinduphobic tweets, declines in sentiment polarity scores, and significant pandemic-related events, such as vaccine distribution controversies and public health debates. Unlike other studies that observed empathetic sentiments in COVID-19-related discourse, Hinduphobic tweets exhibited a stark absence of empathy. This highlights how political narratives and media representations during the pandemic perpetuated communal blame and reinforced harmful stereotypes. Addressing these biases through education, media literacy, and inclusive representation is essential to mitigating communal tensions and fostering a more respectful and informed global discourse.
Data and Code Availability
The open-source code for this study is available on our GitHub repository  [^12] https://github.com/pinglainstitute/Hinduphobia-COVID-19-beyond.
HP-BERT Model Download Link
The HP-BERT model can be downloaded from Zenodo  [^13] https://zenodo.org/records/14607359.
Appendix
Sentiments Sample Tweet Country Score
Optimistic ”Hindu Temples have donated huge sums of money so that people get food and all this is good.” India 0.2
Optimistic, Annoyed ”Annual Hindu pilgrimage Amarnath Yatra has been cancelled this year amid Coronavirus Crisis. But not even a single Hindu is angry at this decision. We are in ’Majority’ but we never misuse it. That’s the beauty of Hinduism!” India 0.2
Annoyed ”Hindu It Cell will be donating 750 N95 Masks to police personnel and frontline workers to help them fight against COVID-19.” India 0.1
Annoyed, Joking ”We can’t call ISIS attacks ’Islamic terrorism’ since it might stir hate against Muslims. We can’t call COVID ’China Virus’ since it might stir hate against Asians. We can’t post crime stats. But we can and must blame all of society’s problems on Hindus.” India 0.0
Optimistic, Joking ”Cow-piss-drinker Hindus made a vaccine to cure COVID. And saved the life of camel-piss drinkers.” UK 0.4
Optimistic, Thankful, Official Report ”The Hindu Religious & Charitable Endowment Dept of Tamil Nadu state government has directed 47 temples to give Rs. 10 crore of surplus income to the CM relief fund.” India 0.5
Optimistic ”Self-quarantine or Yogic turning inwards - How to live with yourself and be happy and healthy.” Brazil 0.3
Optimistic, Thankful, Empathetic ”Dwarkadish Mandir is opened for devotees from today. With proper social distancing, compulsory wearing of masks, use of hand sanitizers, and thermal screening. Welcome all. Jay Dwarkadish.” India 0.5
Annoyed, Denial ”Hindu way of life is pandemic-proof: - Vegetarianism - Social distancing - Frequent bathing and washing - Strict rules on cooking and consuming food - Boost immunity through exercise and stress reduction via yoga and meditation.” India 0.3
Pessimistic, Official Report ”#UK’s Durga Puja and Diwali celebrations turn virtual amid COVID-19.” UK 0.1
Official Report ”5 Hindu Traditions to Help Reduce Coronavirus: Namaste, Vegetarianism, Turmeric and other spices, Pranayama, and Cremation. https://detechter.com/5-hindu-traditions-to-help-reduce-coronavirus/” UK 0.0
Joking ”Indians are so dumb they think cow piss can cure COVID lol, now please take our experimental vaccine.” Japan 0.1
Joking ”Bro, no. Hinduism is the most dangerous thing on the planet. They burnt women with their dead husband. They kill anyone who eats beef. In India, rapes are rampant. They drank cow’s piss and believed it would cure COVID. Their God had an elephant son, three heads, and skin like Avatar.” Australia 0.1
Table 9: Sentiment analysis of Hinduphobic discourse, showcasing positive polarity scores for a diverse sample of tweets from different countries.
Sentiments Sample Tweet Country Score
Annoyed ”Broken England. Tories at England’s bizarrely named Home Office are winding these idiots up to blame refugees, Scots, French, EU, COVID, Blacks, gays, Muslims, Hindus, Sikhs, Russians, Joe Biden, Nicola Sturgeon, but not Tories.” UK -0.1
Annoyed ”Typical western narrative. Blame #Hindus for poor Hindu rate of growth by Antihindu sickular Nehru socialism. Now ensure Hindu Modi doesn’t get credit for #Hindutva mascot #Modi for #NewIndia. Yes, airports, roads, ports, AIIMS, COVID vaccine, and free rations were because of the West.” India -0.1
critical, official report ”RT @NewsHandle: Religious events like Kumbh are reckless and spreading the virus. Why does India allow this?” India -0.3
Pessimistic, Annoyed, Joking ”Nazis were not just mere ideologues; they were the smartest people of their time. In fact, the best scientific minds of the last century were somehow related to or affected by Nazi Germany. Whereas the Hindu scientists are curing COVID with cow piss. Let that sink in…” UK -0.3
Denial, Official Report ”Babaji ka thullu for Baba Ramdev’s Coronil, which was claimed by him to be a foolproof deterrent for the COVID virus: nearly a hundred employees of Ramdev’s Patanjali tested POSITIVE for Coronavirus.” India -0.4
Annoyed ”In view of the COVID-19 pandemic, there are restrictions on large religious congregations. So, no permission has been given to Puja Mandaps to celebrate Ganesh Puja at the community level. Even the Ganapati festival in Mumbai isn’t taking place this year. All are requested to cooperate.” India -0.1
Joking ”Haha no offence, but the last time I heard the whole world laughing was when a MF claimed that cow piss is a cure for COVID-19. Yet still, the world also laughs when people worship cows and drink cow piss. (Never in the future make fun of other religions being a Hindu).” Brazil -0.1
Annoyed, Official Report ”Here comes the Al-Jazeera historian to tarnish 1.5B Hindus. #Hindus run the biggest free food/meal distribution worldwide. Go to any temple in North America, and you will be served a meal. On the contrary, poor Hindus in Bangladesh were asked to chant Islamic verses during floods to receive relief.” India -0.1
Joking ”COVID killed so many in India that they didn’t have enough wood to burn their dead. People started leaving their dead in the rivers. So, are you saying the COVID problem fixed itself? Or was it the cow piss that saved them?” Japan -0.1
Table 10: Sentiment analysis of Hinduphobic discourse, showcasing negative polarity scores for a diverse sample of tweets from different countries.